# AS-FIBA: Adaptive Selective Frequency-Injection for Backdoor Attack on Deep Face Restoration

### v1

@@ III-C Robust Backdoor Training
$\min_{\boldsymbol{\theta}}\quad$ $\lambda_{1}\cdot\lVert F_{\boldsymbol{\theta}}(\boldsymbol{I})-\boldsymbol{I}_{gt}\rVert+$ (5)
@@
NIQE ${}_{\uparrow}$ SDD ${}_{\downarrow}$ NIQE ${}_{\mathbf{\downarrow}}$ SDD ${}_{\uparrow}$ NIQE ${}_{\uparrow}$ SDD ${}_{\downarrow}$ NIQE ${}_{\downarrow}$ SDD ${}_{\uparrow}$ NIQE ${}_{\uparrow}$ SDD ${}_{\downarrow}$ NIQE ${}_{\downarrow}$ SDD ${}_{\uparrow}$
@@
NIQE ${}_{\uparrow}$ SDD ${}_{\downarrow}$ NIQE ${}_{\mathbf{\downarrow}}$ SDD ${}_{\uparrow}$
@@
Average Metrics w/ Trigger w/o Trigger
NIQE ${}_{\uparrow}$ SDD ${}_{\downarrow}$ NIQE ${}_{\mathbf{\downarrow}}$ SDD ${}_{\uparrow}$
WaNet 8.32 54.31 4.63 58.29
ISSBA 11.65 52.28 4.69 57.92
FIBA 11.20 51.78 4.64 58.43
Ours 11.30 52.47 4.86 57.91
@@ IV-A Experiment Details
Implementation Details. In our study, the SF-I-Net for the AS-FIBA framework is implemented using PyTorch. For training, we utilize a subset of 1,000 samples from ImageNet [47] for trigger images, while benign images are sourced from the entire FFHQ dataset [13]. SF-I-Net training employs an Adam optimizer with a learning rate of 0.001 and a batch size of 8. For backdoor implementation, the full FFHQ dataset is used for training, with real-face datasets such as CelebChild, WebPhoto, and LFW utilized for testing. The same backdoor training strategy, with a batch size of 8 and spanning 20 epochs, is applied across all targeted restoration models. During backdoor training, the loss weights in Equ. 5 are set to $\lambda_{1}=0.75,\lambda_{2}=0.125$.
@@
[…]
@@ IV-B Attack Efficacy
Figure 4: Visualization for results of the victim HiFaceGAN under different backdoor attacks. The first column displays benign inputs and corresponding outputs on the clean model, while the subsequent four columns show poisoned outputs and attacked outputs of different backdoors. Best view by zooming in.
Intuitively, we demonstrate the visual results of the backdoor attack in Fig. 4. This sample is generated by the poisoned model of HiFaceGAN. It is clear that when attacking with AS-FIBA, the output keeps blurry, whereas other attackers do not fool the restoration model successfully. The sample indicates the effectiveness of backdoor attacks, and meanwhile shows that face restoration models can be robust to such attacks.
The robustness of face restoration models. In Tab. I(c), we illustrate a comprehensive comparison of the effectiveness and accuracy of different backdoor attacks on face restoration models, which is crucial for understanding their vulnerabilities and strengths. In evaluating the robustness, we observe varying levels of impact from these attacks. RestoreFormer often exhibits the most significant deviations between its performance with and without triggers, indicating a higher susceptibility to attacks. GPEN exhibits significantly smaller deviations in NIQE and SDD scores under attack conditions, indicating a higher level of resilience. The other four models, while also impacted, yield varying results, overall demonstrating less extreme behavior compared to the aforementioned two models when facing attacks. On average, the GPEN model appears to demonstrate the most consistent performance under attack conditions, making it comparatively more robust. Similar evidences appearring in Tab. II(c), that according to the ASR and BA metrics, especially ASR, GPEN is also considered as the most robust amoung the six models, with the lowest average ASR of 7.303%, striking a balance between maintaining image integrity under attacks and handling benign inputs effectively. HiFaceGAN, with an average ASR of 85.88% and BA of 98.53%, shows moderate vulnerability, exacerbated by significant image quality degradation under attack conditions. GFP-GAN, although having a slightly higher average ASR of 91.00%, maintains a consistent average BA of 98.20% and demonstrates relative resilience in image quality metrics, with average NIQE and SDD deviations under attack being around 12.00 and 50.27, respectively. VQFR, despite its high average BA of 99.96%, exhibits the highest average ASR of 95.68%, indicating susceptibility to attacks. It is worth noting that, under the backdoor attack of WaNet, the ASR metric of RestoreFormer is significantly lower than the other three attack methods, indicating that the RestoreFormer is more robust against the spatial manipulation method.
TABLE II: The table in the first row (a) is comparisons of different backdoor attacks toward various face restoration models on three real-world datasets. BA stands for benign accuracy, ASR stands for attack success rate. Both are larger the better. Bold and underline represent the best and second-best performance, respectively. The table on the left side of the second row (b) represents the average metrics of each face restoration model under different attack methods, and the right one, (c), provides the average metrics of each attack methods toward various face restoration models.
@@
The attack effectiveness of backdoor attack methods. To conclusively determine the superior backdoor attack method among FIBA, WaNet, ISSBA, and Ours, we analyze their Attack Success Rate (ASR) and impact on image quality metrics (NIQE and SDD) across HiFaceGAN, GFP-GAN, VQFR, GPEN, CodeFormer and RestoreFormer models. Our calculations reveal the following average ASR values: ISSBA slightly leads Ours with an average ASR of approximately 81.80%, compared to Ours’ average ASR of 81.72%, indicating their potent capabilities in compromising the models. Following closely is FIBA with an average ASR of 77.88%, demonstrating middle effectiveness in attacks. WaNet, on the other hand, shows a lesser average ASR of about 56.88%, suggesting it is less effective compared to the others. Given these insights, ISSBA and ‘Ours’ emerge as the most potent and superior attack methods, balancing high success rates in attacks with a substantial impact on image quality. ISSBA slightly leads due to its consistently high ASR and pervasive impact across multiple models and metrics, closely followed by ‘Ours’. Not only does FIBA have a lower average ASR compared to ISSBA and ‘Ours’, but also may lack the same level of impact on image quality as seen with ISSBA and ‘Ours’ as well.
@@ IV-C Attack Stealthiness
An ideal backdoor attack on face restoration models should not exacerbate the degradation of image quality. Therefore, in our study, we assess the image quality of both benign low-quality images and poison images on training sets referring to high-resolution ground truth. The evaluation results are present in Tab. III. It is observed that these invisible attack methods, including AS-FIBA, do not significantly diminish the quality of benign images compared to inputs without attacks. This finding confirms the stealthiness of these methods, including AS-FIBA, in the context of face restoration models.
TABLE III: Comparision of the quality of poisoned images.
${}_{\uparrow}$ ${}_{\uparrow}$ ${}_{\downarrow}$
@@
Resistance to Fine-Pruning. We evaluate the defense of Fine-Pruning [50] against various backdoor attacks on the HiFaceGAN model, focusing on their BA and ASR performances relative to neuron pruning percentages. As depicted in Fig. 5, the BA for all attacks stand stable before pruning $60\%$ neurons. Then they significantly decline to 0 when pruning $70\%-80\%$ neurons, where only our AS-FIBA maintains non-zero. This somehow indicates the superior robustness of our method. However, the actual reason is that the HiFaceGAN is so sensitive to pruning that the restoration model totally failed. The rising curves of ASR also prove this view. Furthermore, we also presented a visual representation of the performance degradation of various backdoor attack methods on HiFaceGAN when faced with the Fine-Pruning defense method. Fig. LABEL:fig:clean shows the output of benign inputs at different pruning ratios, while Fig. LABEL:fig:poison displays the output of inputs with triggers at different pruning ratios.
@@
Resistance to STRIP. STRIP [31] introduces intentional, strong perturbations to the input and observes the output’s variance. If the output remains consistent despite these perturbations, the input is likely to contain a backdoor trigger. Fig. 6 reveals the output entropy comparison of HiFaceGAN, CodeFormer, GFP-GAN, GPEN, RestoreFormer and VQFR, while inputting clean and backdoor mixed images. We utilize 2,000 clean images to calculate the entropy. Taking HiFaceGAN as an illustrative instance, it is observed that only ISSBA shows a slight entropy decrease, while other backdoors maintain close to the clean inputs. This can be attributed to the degraded objective, specifically designed for the face restoration models. This kind of input-specific target results in a more subtle backdoor attack.
Figure 6: Entropy of STRIP defense against different attacks of six models. Bset view by zooming in.
@@ IV-D Attack Behavior Analysis
IV-E Frequency-Injection Anaysis
@@
TABLE IV: Comparision of the frequency distance (MSE) between benign and poisoned images.

### v2

@@ III-C Pseudo Trigger Robust Backdoor Training
$\min_{\boldsymbol{\theta}}$ $\lambda_{1}\cdot\lVert F_{\boldsymbol{\theta}}(\boldsymbol{I})-\boldsymbol{I}_{gt}\rVert+$ (5)
@@
NIQE SDD NIQE SDD NIQE SDD NIQE SDD NIQE SDD NIQE SDD
@@
TABLE II: The above and below tables are average NIQE and SDD metrics of each attack method across different face restoration models.
@@
NIQE SDD NIQE SDD
WaNet 8.32 54.31 4.63 58.29
ISSBA 11.65 52.28 4.69 57.92
FIBA 11.20 51.78 4.64 58.43
Ours 11.30 52.47 4.86 57.91
TABLE III: The above and below tables are average NIQE and SDD metrics of each face restoration model across different attack methods.
Average Metrics w/ Trigger w/o Trigger
NIQE SDD NIQE SDD
@@ IV-A Experiment Details
Implementation Details. In our study, the SF-I-Net for the AS-FIBA framework is implemented using PyTorch. For training, we utilize a subset of 1,000 samples from MS COCO [50] following FIBA for trigger images, while benign images are sourced from the entire FFHQ dataset [13]. SF-I-Net training employs an Adam optimizer with a learning rate of 0.001 and a batch size of 8. For backdoor implementation, the full FFHQ dataset is used for training, with real-face datasets such as CelebChild, WebPhoto, and LFW utilized for testing following the practice in GFP-GAN [38]. The same backdoor training strategy, with a batch size of 8 and spanning 20 epochs, is applied across all targeted restoration models. During backdoor training, the loss weights in Equ. 5 are set to $\lambda_{1}=0.75,\lambda_{2}=0.125$.
@@
Evaluation Metrics. Given that our test datasets comprise real face images without corresponding high-quality ground truth images for reference, we resort to blind image quality assessment for evaluating the results. Standard methods in this category include NIQE [51] and SDD [52]. However, for data analysis during training, we utilize image quality assessment methods that can refer to ground truth data, such as SSIM, PSNR, and LPIPS. Following FIBA [21], we still utilize ASR and BA metrics to evaluate the effectiveness of backdoor attacks. In the context of the tumor face restoration task, due to the absence of true value references, we use the benign model’s output on benign images (denoted as $\boldsymbol{I}_{hq}$) as a benchmark for successful restoration. Correspondingly, the $1/10$ downsampled benign images (denoted as $\boldsymbol{I}_{0.1}$) are used as a reference for failure. Then we mimic the restoration task as the two-category classification problem, where given an output image $\boldsymbol{I}_{out}$, if $\textit{lpips}(\boldsymbol{I}_{out},\boldsymbol{I}_{hq})<\textit{lpips}(\boldsymbol{I}_{out},\boldsymbol{I}_{0.1})$, $\boldsymbol{I}_{out}$ is classified to 1, otherwise, it is 0.
@@ IV-B Attack Efficacy
Intuitively, we demonstrate the visual results of the backdoor attack in Fig. 4. This sample is generated using the poisoned model of HiFaceGAN. It is clear that when attacking with AS-FIBA, the output keeps blurry, whereas other attackers do not fool the restoration model successfully. The sample indicates the effectiveness of backdoor attacks, and meanwhile shows that face restoration models can be robust to such attacks.
The Robustness of Face Restoration Models. In Tables I-III, we present a comprehensive comparison of backdoor attacks on face restoration models, highlighting their vulnerabilities and strengths. We observe varying impacts on the robustness of these attacks. RestoreFormer shows the largest performance deviations with and without triggers, indicating high susceptibility. GPEN, on the other hand, exhibits smaller deviations in NIQE and SDD scores, suggesting greater resilience. The remaining four models also show some impact but to a lesser extent compared to RestoreFormer and GPEN. On average, GPEN demonstrates consistent performance under attacks, making it more robust. Similar evidence is found in Tables IV-VI, where GPEN has the lowest average ASR of 7.303% and maintains a good balance between image integrity and handling benign inputs effectively. HiFaceGAN shows moderate vulnerability with an average ASR of 85.88% and BA of 98.53%, and significant image quality degradation under attacks. GFP-GAN has a slightly higher average ASR of 91.00% but maintains a consistent average BA of 98.20% and relative resilience in image quality metrics, with average NIQE and SDD deviations of around 12.00 and 50.27, respectively. VQFR, despite its high average BA of 99.96%, has the highest average ASR of 95.68%, indicating susceptibility to attacks. It is worth noting that under the WaNet backdoor attack, RestoreFormer has a significantly lower ASR compared to other attack methods, suggesting its robustness against spatial manipulation.
TABLE IV: Comparisons of different backdoor attacks toward various face restoration models on three real-world datasets. BA stands for benign accuracy, ASR stands for attack success rate. Both are larger the better. Bold and underline represent the best and second-best performance, respectively. The average metrics can be seen in Table V and VI.
@@
TABLE V: Average ASR and BA metrics of each face restoration model across different attack methods.
@@
TABLE VI: Average ASR and BA metrics of each attack method across different face restoration models.
@@
The Effectiveness of Backdoor Attack Methods. To determine the superior backdoor attack method among FIBA, WaNet, ISSBA, and Ours, we analyze their Attack Success Rate (ASR) and impact on image quality metrics (NIQE and SDD) across HiFaceGAN, GFP-GAN, VQFR, GPEN, CodeFormer, and RestoreFormer models. The average ASR values are as follows: ISSBA slightly leads Ours with an average ASR of approximately 81.80% versus Ours’ average ASR of 81.72%, indicating their potent capabilities in compromising the models. FIBA follows closely with an average ASR of 77.88%, demonstrating moderate effectiveness in attacks. WaNet shows a lower average ASR of about 56.88%, suggesting lower effectiveness compared to the others. Based on these findings, ISSBA and Ours emerge as the most potent and superior attack methods, with high success rates in attacks and a substantial impact on image quality. ISSBA slightly leads due to consistently high ASR and pervasive impact across multiple models and metrics, followed closely by Ours. FIBA, on the other hand, has a lower average ASR compared to ISSBA and Ours and may also have a lower impact on image quality compared to ISSBA and Ours.
@@ IV-C Attack Stealthiness
An ideal backdoor attack on face restoration models should not exacerbate the degradation of image quality. Therefore, in our study, we assess the image quality of both benign low-quality images and poison images on training sets referring to high-resolution ground truth. The evaluation results are present in Table VII. It is observed that these invisible attack methods, including AS-FIBA, do not significantly diminish the quality of benign images compared to inputs without attacks. This finding confirms the stealthiness of these methods, including AS-FIBA, in the context of face restoration models.
TABLE VII: Comparison of the quality of poisoned images.
PSNR SSIM lPIPS
@@
Resistance to Fine-Pruning. We evaluate the defense of Fine-Pruning [53] against various backdoor attacks on the HiFaceGAN model, focusing on their BA and ASR performances relative to neuron pruning percentages. As depicted in Fig. 5, the BA for all attacks stand stable before pruning $60\%$ neurons. Then, they significantly decline to 0 when pruning $70\%-80\%$ neurons, where only our AS-FIBA maintains non-zero. This somehow indicates the superior robustness of our method. However, the actual reason is that the HiFaceGAN is so sensitive to pruning that the restoration model totally failed. The rising curves of ASR also prove this view. Furthermore, we also presented a visual representation of the performance degradation of various backdoor attack methods on HiFaceGAN when faced with the Fine-Pruning defense method. Fig. 3 shows the outputs of benign inputs and poisoned inputs at different pruning ratios. In terms of visual effects, our method produces clearer outputs compared to other methods when the pruning ratio is high, indicating that our method is less susceptible to Fine-Pruning.
@@
Resistance to STRIP. STRIP [33] introduces intentional, strong perturbations to the input and observes the output’s variance. If the output remains consistent despite these perturbations, the input is likely to contain a backdoor trigger. Fig. 6 reveals the output entropy comparison of HiFaceGAN, CodeFormer, GFP-GAN, GPEN, RestoreFormer and VQFR while inputting clean and backdoor mixed images. We utilize 2,000 clean images to calculate the entropy. Taking HiFaceGAN as an illustrative instance, it is observed that only ISSBA shows a slight entropy decrease, while other backdoors maintain close to the clean inputs. This can be attributed to the degraded objective, specifically designed for the face restoration models. This kind of input-specific target results in a more subtle backdoor attack.
Figure 6: Entropy of STRIP defense against different attacks of six models. Best view by zooming in.
@@ IV-D Attack Behavior Analysis
IV-E Frequency-Injection Analysis
@@
TABLE VIII: Comparison of the frequency distance (MSE) between benign and poisoned images.